# Axion Icebergs: Clockwork ALPs at hadron colliders

### v1

@@ 5 Vector-like quarks and heavy scalars
6 Summary and Conclusion
We have examined the minimal QCD axion model within the clockwork paradigm and investigated the prospects of observing the massive ALPs that the model engenders at current and future hadron colliders. To this end, we employ a KSVZ-like setup with a heavy $SU(2)_{L}$ singlet quark (as a top-partner) which couples to the CW sector at the site $j=N$ through an analogue of the PQ symmetry and naturally addresses the strong CP problem in the traditional manner. Thus, the lightest pNGB in the CW sector identifies as the QCD axion while the accompanying heavier pNGBs in the spectrum behave as ALPs with a characteristic mass scale $\sim mq$. Through the chiral anomaly, these axions then couple to the gluons, the photon and the $Z$ boson, albeit with hierarchically different decay constants for the QCD axion and the ALPs, namely $f_{0}\sim q^{-N}f$ and $f_{n>0}\sim f$, thanks to the CW mechanism. Being naturally consistent with the current experimental and observational constraints, the large suppression in the effective couplings of the light axion renders it practically invisible at high energy colliders. Therefore, for a reasonable parameter configuration such as $q>1$, $N\sim\mathcal{O}(10)$, $m\gtrsim 10$ GeV and $f\sim 1$ TeV, what seems more plausible instead is the production and detection of the ALPs, especially at hadron colliders such as the currently operating LHC and its future derivatives. To perform a quantitative analysis, we classify three benchmark scenarios according to the ALP masses in the range $10-30$ GeV (BP-I), $35-105$ GeV (BP-II) and $150-450$ GeV (BP-III). Such varied mass scales for the ALPs is enabled by the fact that the CW mechanism is based on the premise of localization in the theory space and, thus, is practically independent of the mass scales assumed in the theory. We performed an analysis of the expected signal profiles pertaining to the three benchmark points for the production of ALPs (with upto two additional jets) via gluon fusion and their subsequent decay to two photons at the LHC for an integrated luminosity of $138{\rm\,fb}^{-1}$. As anchor points for the analysis and for estimates of the pertinent background distributions, we referred to the ATLAS searches for diphoton resonances in the relevant invariant mass regions ATLAS:2022abz ; ATLAS:2021uiz . What stands out as particularly interesting is the case of the light ALPs where the number of particles $N$ that is consistent with the QCD axion being a dark matter candidate also induces very small mass-splittings among the ALPs. The resulting diphoton invariant mass distribution for the signal events turns out to be such that the individual resonances are smeared by the detector’s resolution (implemented in the form of a fast simulation of the ATLAS detector) and, thereby, overlap with each other to a certain extent. Therefore, such scenarios suggest that the existence of a CW spectrum of light ALPs might, in principle, also manifest in the observation of a single broad resonance instead of the typical signatures of distinct, well-separated, resonances. Treating the spectrum as a whole also offers the possibility of a better, more inclusive, estimation of the signal’s statistical significance over the background. For instance, the simplistic cumulative estimation defined in this work gives a sizable significance value of $S_{c}\sim 3.4\sigma$ in BP-I, even though the significance estimates of the individual bins in the signal distribution are comparatively small. Moreover, when scaled by the projected luminosity reach of the ongoing Run 3 phase of the LHC, viz. $\sim 300{\rm\,fb}^{-1}$, the cumulative significance gets enhanced to $S_{c}\sim 5\sigma$. Of course, a dedicated search for such signals would call for a more detailed statistical analysis that is tailored to the fact that the individual resonances of the CW spectrum are characteristically correlated. As for the heavy ALPs, e.g. those encountered in BP-III, one may not be very hopeful within the LHC era, with the cumulative significance falling below $3\sigma$ even at the $300{\rm\,fb}^{-1}$ level. What is exciting, though, is the prospect of probing a wide window of the CW ALPs at the HL-LHC with a planned luminosity reach of $\sim 3000{\rm\,fb}^{-1}$, where larger portions of the diphoton invariant mass profile could start to become apparent — the structure of the axion iceberg revealing itself — and even scenarios like BP-III could possibly surpass the discovery threshold.
@@ 6 Summary and Conclusion
A key issue in our analysis has been the opposing pull on the SSB scale $f$ from the need to increase the effective ALP–gluon–gluon coupling (scaling as $f^{-1}$) on the one hand and the lower bound on the VLQ mass (scaling as $f$) on the other. While modifying the number of sites $N$, or the exponent parameter $q$ does make a difference, the effect is a muted one (especially given that renormalizability of the low energy theory requires that $q\leq 3$). However, given that the light ALP solutions presented here (benchmark points I & II) indicate that further new physics must exist by $\sim 10$ TeV scale, such a restriction might seem unwarranted and $q>3$ may be considered. However, since the ALP couplings to the gluons and photon have a dependence on $q$ which appears only through the elements of the CW transformation matrix, namely $C_{(n>0)N}$, the corresponding gain is marginal. It can be similarly argued that a change in the position of the VLQ coupling over the CW lattice (i.e., to assume $j<N$) also does not lead to a substantial gain.
@@
Naively, a second possibility would be to postulate additional VLQs such that the effective $a_{n}gg$ coupling may be raised by way of all the VLQs contributing in the loop diagram. However, the corresponding change in the axion potential would putatively lead to the formation of stable domain walls in the early Universe and is phenomenologically disfavoured. While one might attempt to mitigate the issue by invoking some of the solutions proposed in the literature, e.g. by assuming a soft breaking of the discrete symmetry or an inflation-induced statistical bias in the axion potential, the model would, nevertheless, be severely constrained due to the inherent drawbacks of such solutions Beyer:2022ywc . Consider, instead, the introduction of a one (or more) vector-like lepton (VLL). This, then, does not contribute to the $a_{n}gg$ couplings (thereby evading the domain wall constraints), but does contribute to the $a_{n}\gamma\gamma$ couplings, which now get enhanced by a factor of
@@
where $N_{L}$ is the number of VLLs of charge $Q_{L}$. Even the simplest choice of a single VLL of charge $Q=-1$ enhances the coupling by a factor of $1.75$, thereby allowing on a similar enhancement in $f$ without changing the rates. This, in turn, allows for a smaller Yukawa $\lambda_{\Psi}$, and postponement of the strong coupling phase until a cutoff scale $\mu\gtrsim 10^{7}$ GeV. As for the VLL phenomenology, the LHC limits are understandably much weaker. For it to decay, it must have at least a small Yukawa coupling with a SM lepton, and postulating this to be the $\tau$ would not only escape low-energy constraints, but also allow for intriguing signals at the LHC.
@@
Whatever such mechanisms may be, it is quite apparent that the clockwork-axion-ALP scenario would be manifested at the LHC in terms of interesting signals, perhaps the most interesting being the case of the strong coupling phase itself.
@@ Appendix A The axial anomaly
$-\frac{\lambda_{\Psi}}{\sqrt{2}}(\phi+f)e^{i\xi\pi_{N}\gamma_{5}/f}\bar{\Psi}\Psi.$ (27)
@@
$\mathcal{D}\bar{\Psi}\mathcal{D}\Psi\to|\mathcal{J}|^{-2}\mathcal{D}\bar{\Psi}\mathcal{D}\Psi\,,$ (29)
where $\mathcal{J}$ is the Jacobian corresponding to the chiral transformation. The Jacobian, after appropriate regularisation, is given by Schwartz_2013 ; Bilal:2008qx
$\mathcal{J}=\frac{i}{64\pi^{2}}\exp{\left[\int d^{4}x\,\frac{\xi\,\beta}{f}\,\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}\right]}.$ (30)
@@
$\mathcal{A}(x)=-\frac{\xi\,\beta}{32\pi^{2}f}\,\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}.$ (31)
@@
$\mathcal{A}(x)=-\frac{\xi\,\pi_{N}}{32\pi^{2}f}\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}.$ (33)
@@ Appendix B VLQ decay widths
$\begin{split}\Gamma(T\to qV)=&|g^{V}_{L}|^{2}\;\frac{\sqrt{m_{q}^{4}-2m_{q}^{2}\left(m_{T}^{2}+m_{V}^{2}\right)+\left(m_{T}^{2}-m_{V}^{2}\right)^{2}}}{32\pi\,m_{T}^{3}\,m_{V}^{2}}\\ &\hskip 30.00005pt\Big{[}\left(m_{T}^{2}-m_{q}^{2}\right)^{2}+m_{V}^{2}\left(m_{T}^{2}+m_{q}^{2}\right)-2m_{V}^{4}\Big{]}\\[6.45831pt] \approx&\frac{|g^{V}_{L}|^{2}}{32\pi}\frac{m_{T}^{3}}{m_{V}^{2}}\quad(\mbox{for}\,\,m_{T}\gg m_{q,V})\quad.\end{split}$ (34)
@@
$\begin{split}\Gamma(T\to t\,h)&=\frac{\sqrt{m_{h}^{4}-2m_{h}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{64\pi\,m_{T}^{3}}\\ &\times\left\{\left(|y^{h}_{Tt}|^{2}+|y^{h}_{tT}|^{2}\right)\left[m_{T}^{2}+m_{t}^{2}-m_{h}^{2}\right]+2\left(y^{h*}_{Tt}y^{h}_{tT}+y^{h}_{Tt}\,y^{h*}_{tT}\right)m_{T}m_{t}\right\}\\[6.45831pt] &\approx\frac{\left(|y^{h}_{Tt}|^{2}+|y^{h}_{tT}|^{2}\right)}{64\pi}m_{T}\quad(\mbox{for}\,\,m_{T}\gg m_{t,h})\ ,\end{split}$ (36)
$\begin{split}\Gamma(T\to t\,\phi_{N})&=\frac{\sqrt{m_{\phi_{N}}^{4}-2m_{\phi_{N}}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{64\pi\,m_{T}^{3}}\\ &\times\left\{\left(|y^{\Phi}_{Tt}|^{2}+|y^{\Phi}_{tT}|^{2}\right)\left[m_{T}^{2}+m_{t}^{2}-m_{\phi_{N}}^{2}\right]+2\left(y^{\Phi*}_{Tt}y^{\Phi}_{tT}+y^{\Phi}_{Tt}\,y^{\Phi*}_{tT}\right)m_{T}m_{t}\right\}\\[6.45831pt] &\approx\frac{\left(|y^{\Phi}_{Tt}|^{2}+|y^{\Phi}_{tT}|^{2}\right)}{64\pi}m_{T}\left(1-2\frac{m_{\phi_{N}}^{2}}{m_{T}^{2}}+\frac{m_{\phi_{N}}^{4}}{m_{T}^{4}}\right)\quad(\mbox{for}\,\,m_{T,\phi_{N}}\gg m_{t})\end{split}$ (37)
@@
$\begin{array}[]{rcl}\Gamma(T\to t\,a_{n})&=&\displaystyle\frac{\xi^{2}C_{nN}^{2}\sqrt{m_{n}^{4}-2m_{n}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{32\pi f^{2}\,m_{T}^{3}}\\[6.45831pt] &&\displaystyle\left(|U_{TT}|^{2}|U_{Tt}|^{2}\right)\left[\left(m_{T}^{2}-m_{t}^{2}\right)^{2}-\left(m_{T}^{2}+m_{t}^{2}\right)m_{n}^{2}\right]\\[10.76385pt] &\approx&\displaystyle\frac{\lambda_{\Psi}^{2}\,\xi^{2}\,C_{nN}^{2}|U_{TT}|^{2}|U_{Tt}|^{2}}{32\pi}\,m_{T}\qquad(\mbox{for}\,\,m_{T}\gg m_{t,a_{n}})\quad,\end{array}$ (38)
@@
$\begin{array}[]{rclcrcl}y^{h}_{Tt}&\approx&\displaystyle\lambda_{h}U^{*}_{Tt}V_{tt}\ ,&&y^{h}_{tT}&=&\displaystyle\lambda_{h}U^{*}_{tt}V_{tT}+y_{\Psi}U^{*}_{tt}V_{TT}\ ,\\[4.30554pt] y^{\Phi}_{tT}&\approx&\displaystyle\lambda_{\Psi}U^{*}_{tT}V_{TT}\ ,&&y^{\Phi}_{Tt}&=&\displaystyle\lambda_{\Psi}U^{*}_{TT}V_{Tt}+y^{\prime}_{\Psi}U^{*}_{TT}V_{tt}\ .\end{array}$ (39)
Acknowledgements.
The authors thank Satyaki Bhattacharya for helpful discussions. S.M. acknowledges research Grant No. CRG/2018/004889 of the SERB, India. T.S. acknowledges the support from the Dr. D.S. Kothari Postdoctoral fellowship scheme no. F.4-2/2006 (BSR)/PH/20-21/0163.

### v2

@@ V The Vector-like quarks and the heavy scalars
VI Summary and Conclusion
We have examined the minimal QCD axion model within the clockwork paradigm and investigated the prospects of observing the massive ALPs that the model engenders at current and future hadron colliders. To this end, we employ a KSVZ-like setup with a heavy coloured $SU(2)_{L}$ singlet quark (as a top-partner) which couples to one end of a CW chain of $(N+1)$ complex scalars. Being chirally charged under the site-specific U(1)U(1)—which acts as an analogue of the PQ symmetry—this quark generates the requisite chiral anomaly for addressing the strong CP problem in a natural manner. The lightest pNGB in the CW sector identifies as the QCD axion while the accompanying heavier pNGBs in the spectrum behave as ALPs with a characteristic mass scale $\sim mq$. Through the chiral anomaly, these axions then couple to the gluons, the photon and the $Z$ boson, albeit with hierarchically [^35] The hierarchy is generated dynamically. different decay constants for the QCD axion and the ALPs, namely $f_{0}\sim q^{N}f$ and $f_{n>0}\sim f$, thanks to the CW mechanism. Being naturally consistent with the current experimental and observational constraints, the large suppression in the effective couplings of the light axion renders it practically invisible at high energy colliders (as is the case with all traditional QCD axions).
For a reasonable parameter configuration such as $q>1$, $N\sim\mathcal{O}(10)$, $m\gtrsim 10$ GeV and $f\sim 1$ TeV, what is more attractive instead is the production and detection of the ALPs, especially at the currently operating LHC and its future derivatives. To perform a quantitative analysis, we classify three benchmark scenarios according to the ALP masses in the range 103010-30 GeV (BP-I), 3510535-105 GeV (BP-II) and 150450150-450 GeV (BP-III). Such varied mass scales for the ALPs are enabled by the fact that the CW mechanism is based on the premise of localization in the theory space and, thus, is practically independent of the mass scales assumed in the theory. We performed an analysis of the expected signal profiles pertaining to the three benchmark points for the production of ALPs (with upto two additional jets) via gluon fusion and their subsequent decay to two photons at the LHC for an integrated luminosity of $138{\rm\,fb}^{-1}$. As anchor points for the analysis and for estimates of the pertinent background distributions, we referred to the ATLAS searches for diphoton resonances in the relevant invariant mass regions Aad et al. (2023a, 2021).
What stands out as particularly interesting is the case of the light ALPs where the number of particles $N$ that is consistent with the allowed window for the QCD axion’s decay constant also corresponds to very small mass-splittings among the ALPs. The resulting diphoton invariant mass distribution for the signal events turns out to be such that the individual resonances are smeared by the detector’s resolution (implemented in the form of a fast simulation of the ATLAS detector) and, thereby, overlap with each other with the full CW spectrum of light ALPs manifesting itself, instead, as a single but very broad resonance. It becomes immediately apparent, though, that a single such resonance is entirely untenable, and that such an excess must have resulted from the merging of several individual narrow resonances. On the other hand, treating the spectrum as a whole also offers the possibility of a better, more inclusive, estimation of the signal’s statistical significance over the background. For instance, the simplistic cumulative estimation defined in this work gives a sizable significance value of $S_{c}\sim 3.4\sigma$ (for an integrated luminosity of only $138~{\rm fb}^{-1}$) in BP-I, even though the significance estimates of the individual bins in the signal distribution are comparatively small. Moreover, when scaled by the projected luminosity reach of the ongoing Run 3 phase of the LHC, viz. $\sim 300{\rm\,fb}^{-1}$, the cumulative significance gets enhanced to $S_{c}\sim 5\sigma$. Of course, a dedicated search for such signals would call for a more detailed statistical analysis that is tailored to the fact that the individual resonances of the CW spectrum are characteristically correlated.
As for the heavy ALPs, e.g. those encountered in BP-III, the cumulative significance hovers just below $3\sigma$ even with $300{\rm\,fb}^{-1}$ of data. It should be borne in mind, though, that our analysis and projections are strictly based on the published ATLAS data. It is conceivable that a more refined analysis would serve to improve the sensitivity. What is exciting, though, is the prospect of probing a wide window of the CW ALPs at the HL-LHC with a planned luminosity reach of $\sim 3000{\rm\,fb}^{-1}$, where larger portions of the diphoton invariant mass profile could start to become apparent — the structure of the axion iceberg revealing itself — and even scenarios like BP-III could possibly surpass the discovery threshold. We note that the kind of multi-ALP signals discussed here possess a very broad and distinct quasi-periodic profile which cannot be reproduced in models with a single heavy ALP or in generic multi-scalar scenarios.
@@ VI Summary and Conclusion
A key issue in our analysis has been the opposing pulls on the SSB scale $f$ from the need to increase the effective ALP–gluon–gluon coupling (scaling as $f^{-1}$) on the one hand and the lower bound on the VLQ mass (scaling as $f$) on the other. While modifying the number of sites $N$, or the exponent parameter $q$ does make a difference, the effect is a muted one (especially given that renormalizability of the low energy theory, as well as vacuum stability, requires that $q\leq 3$). However, given that the light ALP solutions presented here (benchmark points I & II) indicate that further new physics must exist by $\sim 10$ TeV scale, such a restriction might seem unwarranted and $q>3$ may be considered. However, since the ALP couplings to the gluons and photon have a dependence on $q$ which appears only through the elements of the CW transformation matrix, namely $C_{(n>0)N}$, the corresponding gain is marginal. It can be similarly argued that a change in the position of the VLQ coupling over the CW lattice (i.e., to assume $j<N$) also does not lead to a substantial gain.
@@
Naively, a second possibility would be to postulate additional VLQs such that the effective $a_{n}gg$ coupling may be raised by way of all the VLQs contributing in the loop diagram. However, the corresponding change in the axion potential would putatively lead to the formation of stable domain walls/cosmic strings in the early Universe and is phenomenologically disfavoured. While one might attempt to mitigate the issue by invoking some of the solutions proposed in the literature, e.g. by assuming a soft breaking of the discrete symmetry or an inflation-induced statistical bias in the axion potential, the model would, nevertheless, be severely constrained due to the inherent drawbacks of such solutions Beyer and Sarkar (2023). Consider, instead, the introduction of one (or more) vector-like leptons (VLLs). This does not contribute to the $a_{n}gg$ couplings (thereby evading the domain wall constraints), but does contribute to the $a_{n}\gamma\gamma$ couplings, which now get enhanced by a factor of
@@
where $N_{L}$ is the number of VLLs of charge $Q_{L}$. Even the simplest choice of a single VLL of charge Q=1Q=-1 enhances the coupling by a factor of $1.75$, thereby allowing for a similar enhancement in $f$ without changing the rates. This, in turn, allows for a smaller Yukawa $\lambda_{\Psi}$, and postponement of the strong coupling phase until a cutoff scale $\mu\gtrsim 10^{7}$ GeV. As for the VLL phenomenology, the LHC limits are understandably much weaker. For it to decay, it must have at least a small Yukawa coupling with a SM lepton, and postulating this to be the $\tau$ would not only escape low-energy constraints, but also allow for intriguing signals at the LHC.
@@
Whatever such mechanisms may be, it is quite apparent that the clockwork-axion-ALP scenario would be manifested at the LHC in terms of interesting signals, perhaps the most interesting being the case of the possibility of a strong coupling regime.
In addition to diphoton channels, several other prospective final states could enhance our search efforts at the collider, offering complementary avenues to test the model. Of particular interest are processes involving the heavy VLQ or the radial scalars which, while lying below the current LHC sensitivities, can be probed at the upcoming and proposed high-intensity and high-energy upgrades. For instance, as previously mentioned, the exotic decay of the VLQ to ALPs in association with a SM quark, e.g. $T\to t+a_{n}(\to\gamma\gamma)$, presents an interesting channel to study. Processes mediated by the heavy radial scalars $\phi_{n}$ such as $pp\to\phi_{n}\to a_{n}(\to 2\gamma/2j)a_{n}(\to 2\gamma/2j)$ are also worth exploring, more so because of the recent interest in the search for boosted diphotons Hayrapetyan et al. (2024). Another noteworthy process is the axion mediated production of a $Z$ boson in association with a photon, with the $Z$ subsequently decaying into dijets $(pp\to a_{n}\to Z(\to jj)\,\gamma)$. While the cross section for this process is low for the chosen benchmark points in the parameter space and the background is substantial, it remains a promising channel for probing the model at future colliders, particularly for benchmark point III.
This apart, there also exist other intriguing channels, the nonstandard decay $Z\to a_{n}\gamma$ being one interesting example. Although the process is hardly relevant for an investigation at the LHC, given the strong limits placed on the $Z$ exotic decay fractions at LEP, it could be well within the projected sensitivities of future lepton colliders, e.g. the proposed ILC running in the Giga-$Z$ mode Steinberg and Wells (2021). Furthermore, while the minimal setup considered in this work does not engender coupling of the ALPs to the light quarks or leptons even at the one-loop level, one could envision extending or modifying the model to include mixing of the VLQ with the light quarks and introduce additional VLLs that mix with the SM leptons in order to induce ALP couplings with the light SM fermions. In that case, several interesting channels open up, most notable of them being $h\to a_{n}a_{n}\to b\bar{b}\,\mu\bar{\mu}$ (with both the ATLAS and CMS experiments having observed intriguing signals Aad et al. (2022); Sirunyan et al. (2019) in this channel) and $h\to Z\,a_{n},\,a_{n}\to\,\mbox{hadrons}$ Schmieden (2021). This would also have interesting ramifications for flavour physics, including tests for lepton flavour universality in $B$-decays.
A complementary test of the model can also be found in the context of the direct detection of the light axion. A quick inspection shows that the masses and the effective photon couplings of the QCD axion in BP I and II(a) are close to the current sensitivities of several cavity haloscope experiments such as ADMX, CAPP, HAYSTAC, QUAX, etc. Thus, a collider probe of the characteristic multi-ALP signatures, which would largely constrain the parameters $m$ and $f_{\rm eff}$, alongwith the complementary probe of the light axion at the current or next generation of direct detection experiments, which is sensitive to the parameters $f_{\rm eff}$, $q$ and $N$, present a robust experimental platform to test the model in the near future.
Acknowledgements.
The authors thank Satyaki Bhattacharya, Tuhin S. Roy and Shilpi Jain for helpful discussions. D.C. acknowledges IoE, University of Delhi grant IoE/2024-25/12/FRP and grant no. CRG/2023/008234 of the ANRF, India. S.M. acknowledges research Grant No. CRG/2018/004889 of the SERB, India. T.S. acknowledges the support from the Dr. D.S. Kothari Postdoctoral fellowship scheme no. F.4-2/2006 (BSR)/PH/20-21/0163.
Appendices
@@ Appendix A The axial anomaly
$\frac{\lambda_{\Psi}}{\sqrt{2}}(\phi+f)e^{i\xi\pi_{N}\gamma_{5}/f}\bar{\Psi}\Psi.$ (28)
@@
$\mathcal{D}\bar{\Psi}\mathcal{D}\Psi\to\mathcal{J}^{-2}\mathcal{D}\bar{\Psi}\mathcal{D}\Psi\,$, (30)
where $\mathcal{J}$ is the Jacobian corresponding to the chiral transformation. The Jacobian, after appropriate regularisation, is given by Schwartz (2013); Bilal (2008)
$\mathcal{J}=\exp{\left[i\int d^{4}x\,\frac{1}{32\pi^{2}}\frac{\xi\,\beta}{f}\,\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}\right]}.$ (31)
@@
$\mathcal{A}(x)=-\frac{\xi\,\beta}{16\pi^{2}f}\,\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}.$ (32)
@@
$\mathcal{A}(x)=-\frac{\xi\,\pi_{N}}{16\pi^{2}f}\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}.$ (34)
@@ Appendix B VLQ decay widths
$\begin{split}\Gamma(T\to qV)=&|g^{V}_{L}|^{2}\;\frac{\sqrt{m_{q}^{4}-2m_{q}^{2}\left(m_{T}^{2}+m_{V}^{2}\right)+\left(m_{T}^{2}-m_{V}^{2}\right)^{2}}}{32\pi\,m_{T}^{3}\,m_{V}^{2}}\Big{[}\left(m_{T}^{2}-m_{q}^{2}\right)^{2}+m_{V}^{2}\left(m_{T}^{2}+m_{q}^{2}\right)-2m_{V}^{4}\Big{]}\\[6.45831pt] \approx&\frac{|g^{V}_{L}|^{2}}{32\pi}\frac{m_{T}^{3}}{m_{V}^{2}}\quad(\mbox{for}\,\,m_{T}\gg m_{q,V})\quad.\end{split}$ (35)
@@
$\begin{split}\Gamma(T\to t\,h)&=\frac{\sqrt{m_{h}^{4}-2m_{h}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{64\pi\,m_{T}^{3}}\\ &\times\left\{\left(|y^{h}_{Tt}|^{2}+|y^{h}_{tT}|^{2}\right)\left[m_{T}^{2}+m_{t}^{2}-m_{h}^{2}\right]+4\mbox{Re}\left(y^{h*}_{Tt}y^{h}_{tT}\right)m_{T}m_{t}\right\}\\[6.45831pt] &\approx\frac{|y^{h}_{Tt}|^{2}+|y^{h}_{tT}|^{2}}{64\pi}m_{T}\quad(\mbox{for}\,\,m_{T}\gg m_{t,h})\ ,\end{split}$ (37a)
$\begin{split}\Gamma(T\to t\,\phi_{N})&=\frac{\sqrt{m_{\phi_{N}}^{4}-2m_{\phi_{N}}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{64\pi\,m_{T}^{3}}\\ &\times\left\{\left(|y^{\Phi}_{Tt}|^{2}+|y^{\Phi}_{tT}|^{2}\right)\left[m_{T}^{2}+m_{t}^{2}-m_{\phi_{N}}^{2}\right]+4\mbox{Re}\left(y^{\Phi*}_{Tt}y^{\Phi}_{tT}\right)m_{T}m_{t}\right\}\\[6.45831pt] &\approx\frac{|y^{\Phi}_{Tt}|^{2}+|y^{\Phi}_{tT}|^{2}}{64\pi}m_{T}\left(1-\frac{m_{\phi_{N}}^{2}}{m_{T}^{2}}\right)^{2}\quad(\mbox{for}\,\,m_{T,\phi_{N}}\gg m_{t})\end{split}$ (37b)
@@
$\begin{array}[]{rcl}\Gamma(T\to t\,a_{n})&=&\displaystyle\frac{\xi^{2}C_{nN}^{2}\sqrt{m_{n}^{4}-2m_{n}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{32\pi f^{2}\,m_{T}^{3}}\\[6.45831pt] &&\displaystyle\left(|U_{TT}|^{2}|U_{Tt}|^{2}\right)\left[\left(m_{T}^{2}-m_{t}^{2}\right)^{2}-\left(m_{T}^{2}+m_{t}^{2}\right)m_{n}^{2}\right]\\[10.76385pt] &\approx&\displaystyle\frac{\lambda_{\Psi}^{2}\,\xi^{2}\,C_{nN}^{2}|U_{TT}|^{2}|U_{Tt}|^{2}}{32\pi}\,m_{T}\qquad(\mbox{for}\,\,m_{T}\gg m_{t,a_{n}})\quad,\end{array}$ (37c)
@@
$\begin{split}y^{h}_{Tt}&\approx\displaystyle\lambda_{h}U^{*}_{Tt}V_{tt}\ ,\\ y^{h}_{tT}&=\displaystyle\lambda_{h}U^{*}_{tt}V_{tT}+y_{\Psi}U^{*}_{tt}V_{TT}\ ,\\ y^{\Phi}_{tT}&\approx\displaystyle\lambda_{\Psi}U^{*}_{tT}V_{TT}\ ,\\ y^{\Phi}_{Tt}&=\displaystyle\lambda_{\Psi}U^{*}_{TT}V_{Tt}+y^{\prime}_{\Psi}U^{*}_{TT}V_{tt}\ .\end{split}$ (38)